### v2

# Shallow Flow Matching for Coarse-to-Fine Text-to-Speech Synthesis

Abstract
We propose Shallow Flow Matching (SFM), a novel mechanism that enhances flow matching (FM)-based text-to-speech (TTS) models within a coarse-to-fine generation paradigm. Unlike conventional FM modules, which use the coarse representations from the weak generator as conditions, SFM constructs intermediate states along the FM paths from these representations. During training, we introduce an orthogonal projection method to adaptively determine the temporal position of these states, and apply a principled construction strategy based on a single-segment piecewise flow. The SFM inference starts from the intermediate state rather than pure noise, thereby focusing computation on the latter stages of the FM paths. We integrate SFM into multiple TTS models with a lightweight SFM head. Experiments demonstrate that SFM yields consistent gains in speech naturalness across both objective and subjective evaluations, and significantly accelerates inference when using adaptive-step ODE solvers. Demo and codes are available at https://ydqmkkx.github.io/SFMDemo/.
1 Introduction
Text-to-speech (TTS) synthesis has advanced with generative algorithms in recent years, particularly autoregressive (AR) VALL-E ; Fish-Speech ; base-tts ; llasa ; llmvox ; spark-tts , diffusion Grad-TTS ; E3 ; InstructTTS ; Seed-TTS ; DiTTo-TTS , and flow matching (FM) Voicebox ; Reflow-TTS ; voiceflow ; e2-tts ; matcha-tts ; f5-tts ; f5r-tts methods. TTS models typically contain a front-end for processing contextual information and a back-end for speech synthesis. Due to the strong temporal alignment between texts and speech, many diffusion- or FM-based TTS models adopt a coarse-to-fine generation paradigm: a weak generator first produces coarse representations conditioned on the input context, which are then refined by the diffusion or FM module into high-quality mel-spectrograms, and finally converted into audio waveforms by a vocoder. There are two main approaches to the weak generator. One Grad-TTS ; Reflow-TTS ; voiceflow ; matcha-tts follows traditional TTS designs, where a non-autoregressive encoder and an alignment module fastspeech2 ; vits jointly generate coarse mel-spectrograms when aligning encoder outputs with target mel-spectrograms. The other Seed-TTS ; cosyvoice ; cosyvoice2 ; melle ; minimaxspeech employs an AR large language model (LLM) as a context processor and weak generator to generate discrete speech tokens, which are transformed into continuous mel-spectrograms by a diffusion or FM module.
FM has gained increasing attention due to its efficiency and high-quality generation in TTS. In conventional coarse-to-fine FM-based TTS models, coarse representations are used as conditions for the flow module. However, the generation still starts from pure noise, resulting in a suboptimal allocation of modeling capacity. Since the coarse representations already encode a significant portion of the overall semantic and acoustic structure, modeling the early stage from noise becomes redundant and contributes little to the quality of the final output. To deal with this issue, DiffSinger diffsinger has introduced a shallow diffusion mechanism on singing voice synthesis (SVS) and TTS. It uses a simple mel-spectrogram decoder, with the diffusion module starting generation at a shallow step based on the output of the decoder. Therefore, we extend the idea of shallow diffusion mechanism (with the name) on FM and propose shallow flow matching (SFM) for coarse-to-fine TTS models. During training, we construct intermediate states along the FM paths based on the coarse representations. To achieve this, we employ an orthogonal projection method to adaptively determine the corresponding time, define a principled construction approach, and formulate a single-segment piecewise flow. During inference, the generation starts from the intermediate state, skipping the early stages and focusing computation on the latter part of the flow. This leads to more stable generation and enhanced synthesis quality, and accelerates inference when using adaptive-step ordinary differential equation (ODE) solvers.
We validate the proposed SFM method on multiple coarse-to-fine TTS models, covering two mainstream FM architectures, U-Net unet and DiT dit . Experimental results show that SFM consistently improves the naturalness of synthesized speech, as measured by both pseudo-MOS metrics and subjective evaluations. We also observe significant inference acceleration across various adaptive-step ODE solvers.
2 Preliminaries
2.1 Flow matching
A time-dependent diffeomorphic map $\boldsymbol{\phi}_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ describes the smooth and invertible transformation of data points $\boldsymbol{x}\in\mathbb{R}^{d}$ over time $t\in[0,1]$, then a flow is defined via the ODE with a time-dependent vector field (VF) $\boldsymbol{u}_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$:
$\boldsymbol{x}_{t}=\boldsymbol{\phi}_{t}(\boldsymbol{x}_{0}),\quad\frac{d}{dt}\boldsymbol{\phi}_{t}(\boldsymbol{x}_{0})=\boldsymbol{u}_{t}(\boldsymbol{\phi}_{t}(\boldsymbol{x}_{0})).$ (1)
[…]
$\boldsymbol{\phi}_{t}(\boldsymbol{x}_{t})=\boldsymbol{\sigma}_{t}(\boldsymbol{x}_{1})\boldsymbol{x}_{t}+\boldsymbol{\mu}_{t}(\boldsymbol{x}_{1}),\quad\boldsymbol{u}_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{1})=\frac{\boldsymbol{\sigma}_{t}^{\prime}(\boldsymbol{x}_{1})}{\boldsymbol{\sigma}_{t}(\boldsymbol{x}_{1})}(\boldsymbol{x}_{t}-\boldsymbol{\mu}_{t}(\boldsymbol{x}_{1}))+\boldsymbol{\mu}_{t}^{\prime}(\boldsymbol{x}_{1}),$ (2)
where $\boldsymbol{\mu}_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the time-dependent mean and $\boldsymbol{\sigma}_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ is the time-dependent scalar standard deviation (std). $f^{\prime}$ denotes the derivative with respect to time, $f^{\prime}=\frac{d}{dt}f$. Corresponding to the optimal transport (OT) displacement interpolant ot , the mean and std are:
$\boldsymbol{\mu}_{t}(\boldsymbol{x}_{1})=t\boldsymbol{x}_{1},\quad\sigma_{t}(\boldsymbol{x}_{1})=1-(1-\sigma_{\mathrm{min}})t,$ (3)
where $\sigma_{\mathrm{min}}$ is a sufficiently small value. Substituting Eq. (3) into Eq. (2), the conditional flow and VF take the form:
$\boldsymbol{\phi}_{t}(\boldsymbol{x}_{0})=(1-t)\boldsymbol{x}_{0}+t(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),\quad\boldsymbol{u}_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{1})=(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})-\boldsymbol{x}_{0}.$ (4)
Then, we can minimize the conditional flow matching (CFM) loss during training, which is proven by fm to be equivalent to minimizing the FM loss $\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}_{t,p_{t}(\boldsymbol{x}_{t})}\|\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)-\boldsymbol{u}_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{1})\|^{2}$.
During inference, we use an ODE solver to solve the integral $\boldsymbol{x}_{\mathrm{pred}}=\boldsymbol{x}_{0}+\int_{0}^{1}\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)\,dt$.
2.2 Classifier-free guidance
In FM-based generative models, a typical approach for controlling the generation process is to incorporate the input condition $\boldsymbol{c}$ during training and inference. To improve diversity and fidelity, classifier-free guidance (CFG) cfg can be employed. During training, the condition $\boldsymbol{c}$ is randomly dropped, enabling the model to learn from both conditional and unconditional contexts. During inference, a hyperparameter called the CFG strength $\beta>0$ is introduced to control the trade-off:
$\boldsymbol{v}_{\boldsymbol{\theta},\mathrm{CFG}}(\boldsymbol{x}_{t},t,\boldsymbol{c})=\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})+\beta(\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t,\boldsymbol{c})-(\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)).$ (5)
Specifically, the FM module runs two forward passes at each time step, once with $\boldsymbol{c}$ and once without.
2.3 Flow matching-based TTS models
Our backbone configurations involve three fully open-source TTS models.
Matcha-TTS: Matcha-TTS matcha-tts is a non-AR FM-based TTS model that employs a conventional encoder-decoder architecture. The Transformer-based attention encoder takes phonemes and speaker IDs (for multi-speaker training) as input, producing hidden states and predicted phoneme-level durations. The U-Net-based FM decoder receives the encoder’s outputs and speaker embeddings as FM conditions and generates mel-spectrograms.
Because the encoder adopts the monotonic alignment search (MAS)-based alignment glow-tts ; vits ; Grad-TTS for phoneme-spectrogram alignment, the encoder outputs are optimized towards the ground-truth mel-spectrograms (via the prior loss in the paper), resulting in coarse mel-spectrograms.
CosyVoice: CosyVoice cosyvoice is a large zero-shot TTS model that consists of four components: the text encoder, the speech tokenizer, the LLM, and the FM module. The speech tokenizer extracts discrete speech tokens from mel-spectrograms of waveforms, while the text encoder processes textual inputs and aligns text encodings with speech tokens. The LLM, a Transformer decoder-based model, takes speaker embeddings, text encodings, and prompt speech tokens as input and generates target speech tokens in an AR way. Subsequently, the FM module, conditioned on speaker embeddings, target speech tokens, and masked mel-spectrograms, generates the target mel-spectrograms.
The FM module adopts an encoder-decoder structure. The Conformer-based conformer encoder encodes the speech tokens generated by the LLM into hidden states and linearly projects them to the same dimension as mel-spectrograms. These hidden states are then upsampled by a length regulator for token-spectrogram alignment and fed into the U-Net-based decoder, along with other conditions, for FM training and inference. Therefore, the upsampled hidden states can be explicitly supervised to coarse mel-spectrograms.
In CosyVoice, the speech tokenizer, LLM, and flow modules are trained independently. Therefore, in our experiments, we only train the flow module using speech tokens generated by the officially pre-trained speech tokenizer, while employing the officially pre-trained LLM during inference.
StableTTS: StableTTS stableTTS is an open-source TTS model in which both the encoder and FM decoder leverage DiT blocks. Its encoder employs MAS-based alignment and outputs coarse mel-spectrograms.
To assess the effectiveness of our method under different input types (text sequences and speech tokens) within the DiT architecture, we further adapt StableTTS as the FM module of CosyVoice. The resulting backbone TTS model is referred to as CosyVoice-DiT throughout our work.
3 Method
3.1 Theorems
Proofs of our theorems are provided in Appendix A. Theorem 2 can be derived from PeRFlow PeRFlow that divides CondOT paths into several windows and conducts piecewise reflow Rectified_Flow in each window.
Theorem 1.
For any random variable $\boldsymbol{x}_{m}\sim\mathcal{N}(t_{m}\boldsymbol{x}_{1},\sigma^{2}_{m}\boldsymbol{I})$, where $t_{m}\in[0,\infty)$ and $\sigma_{m}\in(0,\infty)$, we define a transformation that maps $\boldsymbol{x}_{m}$ onto the conditional OT (CondOT) paths. The output distribution varies continuously with respect to $t_{m}$ and $\sigma_{m}$ under the Wasserstein-2 metric.
$\Delta=(1-\sigma_{\mathrm{min}})t_{m}+\sigma_{m},$ (6)
$\boldsymbol{x}_{\tau}$ $=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}}\boldsymbol{x}_{0}+\boldsymbol{x}_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}\boldsymbol{x}_{m},&\text{if }\Delta\geq 1,\end{cases}$ (7)
where $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, with corresponding $\tau=\min(t_{m},\frac{t_{m}}{\Delta})$ on the path.
Theorem 2.
For arbitrary intermediate states on the CondOT paths:
$\boldsymbol{x}_{t_{m}}=(1-t_{m})\boldsymbol{x}_{0}+t_{m}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),\quad t_{m}\in(0,1),\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}),$ (8)
we can divide the paths into two segments at $t_{m}$ and represent the flow and VF using piecewise functions:
$\boldsymbol{x}_{t}$ $=\begin{cases}(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}},&\text{if }t<t_{m},\\ (1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),&\text{if }t\geq t_{m},\end{cases}$ (9)
$\boldsymbol{u}_{t}$ $=\begin{cases}\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0}),&\text{if }t<t_{m},\\ \frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}}),&\text{if }t\geq t_{m}.\end{cases}$ (10)
3.2 Proposed methods on coarse-to-fine FM-based TTS
Figure 1: Inference process. Left: standard FM; Right: proposed SFM.
We define the mel-spectrogram of an audio waveform as $\boldsymbol{X}\in\mathbb{R}^{N\times F}$, where $N$ is the number of time frames and $F$ is the number of frequency bins (channels). Specifically, $\boldsymbol{X}^{n}\in\mathbb{R}^{F}$ denotes the $n$-th mel-spectrogram frame. We refer to the weak generator as $\boldsymbol{g_{\omega}}$ (with learnable parameters $\boldsymbol{\omega}$), which receives text, speaker features, and other contextual information as the input condition $\boldsymbol{C}$ and outputs a coarse mel-spectrogram $\hat{\boldsymbol{X}}_{\boldsymbol{g}}$. $\hat{\boldsymbol{X}}_{\boldsymbol{g}}$ is supervised to match the target sample $\boldsymbol{X}_{1}$, typically using an L2 loss, though other loss functions may also be applied depending on the task:
[…]
We introduce a lightweight SFM head $\boldsymbol{h_{\psi}}$ with learnable parameters $\boldsymbol{\psi}$ (see details in Appendix B), which takes the final hidden states $\hat{\boldsymbol{H}}_{\boldsymbol{g}}$ from $\boldsymbol{g_{\omega}}$ as input and outputs a scaled mel-spectrogram $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$. Note that $\hat{\boldsymbol{X}}_{\boldsymbol{g}}$ is obtained by applying a linear projection to $\hat{\boldsymbol{H}}_{\boldsymbol{g}}$. In addition, $\boldsymbol{h_{\psi}}$ needs to predict a time $\hat{t}_{\boldsymbol{h}}\in(0,1)$ and an estimated variance $\hat{\sigma}^{2}_{\boldsymbol{h}}$ for $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$:
$\hat{\boldsymbol{H}}_{\boldsymbol{g}},\hat{\boldsymbol{X}}_{\boldsymbol{g}}=\boldsymbol{g_{\omega}}(\boldsymbol{C}),\quad\hat{\boldsymbol{X}}_{\boldsymbol{h}},\hat{t}_{\boldsymbol{h}},\hat{\sigma}^{2}_{\boldsymbol{h}}=\boldsymbol{h_{\psi}}(\hat{\boldsymbol{H}}_{\boldsymbol{g}}).$ (12)
3.2.1 Orthogonal projection onto CondOT paths
Since the exact location of $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ on the CondOT paths and the corresponding time $t_{\boldsymbol{h}}$ are unknown, we let the model adaptively determine $t_{\boldsymbol{h}}$ during training. To direct $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ to the CondOT paths, we find the orthogonal projection of $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ onto $\boldsymbol{X}_{1}$. According to Eq. (3), the projection coefficient is $t_{\boldsymbol{h}}$, which corresponds to an intermediate state on the mean path $\boldsymbol{\mu}_{t}(\boldsymbol{X}_{1})$. Then we estimate the time $t_{\boldsymbol{h}}$ and the variance $\sigma_{\boldsymbol{h}}^{2}$, and minimize the distance between $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ and $t_{\boldsymbol{h}}\boldsymbol{X}_{1}$ via the loss $\mathcal{L}_{\mu}$:
[…]
where $\mathrm{sg}[\cdot]$ (stop gradient) is used to simplify gradient propagation. $\sigma_{\boldsymbol{h}}^{2}$ represents the noise scale of $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ and can be interpreted as intrinsic noise. Given the large number of mel-spectrogram frames, we omit the unbiased correction term in the estimation.
When $\hat{\boldsymbol{X}}_{\boldsymbol{h}}\approx t_{\boldsymbol{h}}\boldsymbol{X}_{1}$, we can assume that $\hat{\boldsymbol{X}}_{\boldsymbol{h}}\sim\mathcal{N}(t_{\boldsymbol{h}}\boldsymbol{X}_{1},\sigma^{2}_{\boldsymbol{h}}\boldsymbol{I})$ and utilize Theorem 1 to construct intermediate states on the CondOT paths:
$\Delta=\max((1-\sigma_{\mathrm{min}})t_{\boldsymbol{h}}+\sigma_{\boldsymbol{h}},1),\quad\tilde{\boldsymbol{X}}_{\boldsymbol{h}}=\frac{1}{\Delta}{\hat{\boldsymbol{X}}_{\boldsymbol{h}}},\quad\tilde{t}_{\boldsymbol{h}}=\frac{1}{\Delta}{t_{\boldsymbol{h}}},\quad\tilde{\sigma}^{2}_{\boldsymbol{h}}=\frac{1}{\Delta^{2}}{\sigma^{2}_{\boldsymbol{h}}},$ (14)
$\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}=\sqrt{\max((1-(1-\sigma_{\mathrm{min}})\tilde{t}_{\boldsymbol{h}})^{2}-\tilde{\sigma}_{\boldsymbol{h}}^{2},0)}\boldsymbol{X}_{0}+\tilde{\boldsymbol{X}}_{\boldsymbol{h}},$ (15)
where $\boldsymbol{X}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. During the early stages of training, it is possible that $\Delta\geq 1$. In such cases, $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ is unable to lie on the CondOT path with the external noise $\boldsymbol{X}_{0}$, resulting in deterministic model behavior. The scaling factor $\frac{1}{\Delta}$ introduced in Theorem 1 rescales and guides $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ back onto the CondOT paths until $\boldsymbol{X}_{0}$ can be incorporated. Then we calculate the losses for the two predicted scalars:
$\mathcal{L}_{t}=(\hat{t}_{\boldsymbol{h}}-\tilde{t}_{\boldsymbol{h}})^{2},\quad\mathcal{L}_{\sigma}=(\hat{\sigma}^{2}_{\boldsymbol{h}}-\tilde{\sigma}^{2}_{\boldsymbol{h}})^{2}.$ (16)
3.2.2 Single-segment piecewise flow
During training and inference, the flow starts from $\tilde{t}_{\boldsymbol{h}}$. Therefore, we employ Theorem 2 and focus only on the second segment of the paths ($t\geq\tilde{t}_{\boldsymbol{h}}$):
$t_{\mathcal{U}}\sim\mathcal{U}[0,1],\quad t_{\mathcal{S}}=\mathcal{S}(t_{\mathcal{U}}),\quad t=(1-\tilde{t}_{\boldsymbol{h}})t_{\mathcal{S}}+\tilde{t}_{\boldsymbol{h}},$ (17)
$\boldsymbol{X}_{t}$ $=\left(1-\frac{t-\tilde{t}_{\boldsymbol{h}}}{1-\tilde{t}_{\boldsymbol{h}}}\right)\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}+\frac{t-\tilde{t}_{\boldsymbol{h}}}{1-\tilde{t}_{\boldsymbol{h}}}(\boldsymbol{X}_{1}+\sigma_{\mathrm{min}}\boldsymbol{X}_{0})$ (18)
$=(1-t_{\mathcal{S}})\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}+t_{\mathcal{S}}(\boldsymbol{X}_{1}+\sigma_{\mathrm{min}}\boldsymbol{X}_{0}),$ (19)
$\boldsymbol{U}_{t}$ $=\frac{1}{1-\tilde{t}_{\boldsymbol{h}}}(\boldsymbol{X}_{1}+\sigma_{\mathrm{min}}\boldsymbol{X}_{0}-\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}),$ (20)
where $\mathcal{S}$ is an arbitrary time scheduler for the randomly sampled $t$.
Finally, combined with the CFM loss, the overall loss for the SFM framework is given by:
[…]
3.2.3 Inference with SFM strength
During inference, we observe that the adaptively determined $t_{\boldsymbol{h}}$ tends to be small, which limits the amount of prior information. To address this issue, we introduce a hyperparameter called SFM strength $\alpha\geq 1$, which scales up the $\hat{t}_{\boldsymbol{h}}$ to encourage stronger guidance from $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$
$\Delta=\max(\alpha((1-\sigma_{\mathrm{min}})\hat{t}_{\boldsymbol{h}}+\hat{\sigma}_{\boldsymbol{h}}),1),\quad\tilde{\boldsymbol{X}}_{\boldsymbol{h}}=\frac{\alpha}{\Delta}{\hat{\boldsymbol{X}}_{\boldsymbol{h}}},\quad\tilde{t}_{\boldsymbol{h}}=\frac{\alpha}{\Delta}{\hat{t}_{\boldsymbol{h}}},\quad\tilde{\sigma}^{2}_{\boldsymbol{h}}=\frac{\alpha^{2}}{\Delta^{2}}\hat{\sigma}^{2}_{\boldsymbol{h}},$ (22)
Substituting $\tilde{\boldsymbol{X}}_{\boldsymbol{h}}$, $\tilde{t}_{\boldsymbol{h}}$, and $\tilde{\sigma}^{2}_{\boldsymbol{h}}$ into Eq. (15) yields $\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}$. We can obtain the predicted results $\boldsymbol{X}_{\mathrm{pred}}$ by solving the integral $\boldsymbol{X}_{\mathrm{pred}}=\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}+\int_{\tilde{t}_{\boldsymbol{h}}}^{1}\boldsymbol{v_{\theta}}(\boldsymbol{X}_{t},t)\,dt$ with an ODE solver. For each model employing the SFM method, we determine the optimal $\alpha$ on the validation set. In most cases, the optimal $\alpha$ is relatively small, resulting in $\Delta=1$. The scaling factor $\frac{\alpha}{\Delta}$ has a theoretical upper bound of $\frac{1}{(1-\sigma_{\mathrm{min}})\hat{t}_{\boldsymbol{h}}+\hat{\sigma}_{\boldsymbol{h}}}$.
We provide concise algorithm boxes in Appendix C, using minimal notation for clarity and including implementation details. An example of SFM inference is illustrated in Fig. 1.
4 Experimental setup
4.1 Datasets
We use LJ Speech ljspeech , VCTK vctk , and LibriTTS libritts in our experiments, where LJ Speech is a single-speaker dataset and the others are multi-speaker datasets. For LJ Speech and VCTK, the training, validation, and test sets are divided following the setting of Matcha-TTS, which follows VITS’s settings. [^1] https://github.com/jaywalnut310/vits/tree/main/filelists Each validation set contains 100 utterances, and each test set contains 500 utterances.
For LibriTTS, the train subsets are used as the training set. We construct the validation and test sets with the dev-clean and test-clean subsets, respectively. We adopt the cross-sentence evaluation approach in CosyVoice and F5-TTS: utterances from each subset are paired to form <prompt, target> pairs, where the prompt utterance is used as the reference to TTS models and the transcript of the target utterance serves as the text input. The prompt utterances are selected such that their durations are rounded to between 3 and 4 seconds, while the duration of target utterances is between 4 and 10 seconds. We set the validation and test sets to contain 200 and 1000 utterances, respectively.
4.2 Model implementations
Since the SFM head (the architecture is shown in Appendix B) receives hidden states as input for richer contextual information, all length regulators in each model upsample the hidden states. Several neural network layers smooth the upsampled hidden states, which are then used to generate coarse mel-spectrograms and input to the SFM head as in Eq. (12). The coarse mel-spectrograms are supervised using the target mel-spectrograms via Eq. (11).
Matcha-TTS: We utilize Matcha-TTS with the officially pre-trained Vocos vocos as the vocoder. Since the official implementation does not include smoothing layers for the encoder outputs, we designate the final block of the six-block encoder to serve as smoothing layers.
StableTTS: Although StableTTS incorporates a reference encoder to extract speaker embeddings and enables zero-shot TTS, we observe that the extracted speaker embeddings are unstable and lead to unstable encoder outputs. This instability causes the simple SFM head to struggle with converging and predicting accurate $\hat{t}_{\boldsymbol{h}}$ and $\hat{\sigma}^{2}_{\boldsymbol{h}}$. Therefore, we use ID-based speaker embeddings with random initialization instead of the reference encoder. Both the encoder and decoder are configured with six DiT blocks. The Vocos pre-trained in StableTTS is used as the vocoder.
CosyVoice: As discussed in Section 2.3, we only train the flow module of CosyVoice. We use the officially pre-trained speech tokenizer and LLM to generate speech tokens for training and inference, respectively. The pre-trained HiFi-GAN hifigan in CosyVoice is used as the vocoder. In the original FM module, the encoder takes only speech tokens as input, while speaker embeddings and masked mel-spectrograms are used as conditions for the FM decoder. However, under this configuration, we find that the encoder fails to produce coarse mel-spectrograms with sufficient speaker features. To address this, we concatenate the speech tokens, speaker embeddings, and masked mel-spectrograms, and feed them jointly into the encoder. For ablation studies, we evaluate a variant (denoted as SFM-t) with the SFM method in which the encoder takes only speech tokens as input.
CosyVoice-DiT: As discussed in Section 2.3, we construct CosyVoice-DiT using the encoder and decoder of StableTTS, each of which contains six DiT blocks. The speech tokens are fed into the encoder, while the speaker embeddings are incorporated into every DiT block via adaLN-Zero dit . The Vocos pre-trained in StableTTS is used as the vocoder.
Baseline models: We adopt and train the original models as baseline models. Due to necessary architectural modifications, no direct baseline is available for StableTTS and CosyVoice-DiT.
Ablated models: For each model, we construct an ablated version to enable a more direct comparison with the SFM model, mainly including the addition of coarse loss and the SFM head. In the ablated models, the SFM head only outputs the $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ as the condition for the FM module. Therefore, the only difference between the ablated and SFM models is whether the SFM method is applied. This design allows us to isolate and assess the effect of the proposed SFM method.
SFM-c: In our SFM method, the coarse mel-spectrograms are not used as FM conditions. For ablation studies, we also evaluate a variant (denoted as SFM-c) that not only applies the SFM method but also uses coarse mel-spectrograms as FM conditions.
4.3 Training and inference configurations
Table 1: Training and inference configurations of TTS models. $\beta$: CFG strength. $\dagger$}: 24,000 maximum frames.
Model Vocoder Dataset Epochs Batch Size Learning Rate Warmup GPUs $\boldsymbol{\beta}$ Solver
[…]
CosyVoice-DiT Vocos LibriTTS 200 256 $\times1$0-4 10% 4 3 Dopri(5)
All training, inference, and objective evaluations are conducted on 96 GB Nvidia H100 GPUs with half precision (FP16). We follow the official configurations of each baseline as closely as possible, and some key settings are summarized in Table 1. Specifically, a warmup parameter indicates that a learning rate scheduler is used. Because of our large batch sizes, we increase the constant or peak learning rates based on the linear scaling rule, and then reduce them if gradient explosions are observed. The presence of a CFG strength indicates the application of CFG.
4.4 Evaluations
4.4.1 Objective evaluation
We evaluate model performance using three pseudo-MOS prediction models: UTMOS utmos , UTMOSv2 utmosv2 , and Distill-MOS distilmos . Their average is defined as $\overline{\text{PMOS}}$ and is used for selecting the optimal $\alpha$ in SFM methods. In addition, we report word error rate (WER) and speaker similarity (SIM). For WER, we use Whisper-large-v3 whisper as the speech recognition model to transcribe speech. For SIM, we use WavLM-base-plus-sv wavlm ; xvector  [^2] https://huggingface.co/microsoft/wavlm-base-plus-sv to extract speaker embeddings and compute the cosine similarity between synthesized and ground-truth speech.
4.4.2 Subjective evaluation
We conducted comparative MOS (CMOS) and similarity MOS (SMOS) tests to evaluate naturalness and similarity, respectively. In each single test, we recruited 20 native English listeners on Prolific [^3] https://www.prolific.com with £9 per hour.
CMOS: We use our proposed model with SFM as the reference system, and all other systems are compared against it. For each system, five utterances are selected and paired with corresponding utterances from the SFM model. Listeners are presented with utterance pairs and asked to rate the naturalness difference on a 7-point scale ranging from 3-3 to +3+3. A score of +3+3 indicates that the first utterance sounds much better than the second, while 3-3 indicates the opposite.
SMOS: SMOS tests are only conducted for cross-sentence evaluations to evaluate the similarity between the ground-truth prompt utterance and the target utterance from different systems. Each system provides five utterances with corresponding prompts. Listeners are presented with <prompt, target> utterance pairs and asked to rate how similar the target sounds to the prompt on a 5-point scale ranging from 1 to 5, where 1 indicates "not similar at all" and 5 indicates "extremely similar".
4.5 SFM strength selection
We determine the optimal SFM strength $\alpha$ on the corresponding validation sets. Starting from 1.0, we increment $\alpha$ by 1.0 and objectively evaluate the generated speech to identify the value that yields the highest $\overline{\text{PMOS}}$. We then examine its neighbors at $\alpha\pm 0.5$ for potential improvements. Finer-grained search is avoided to prevent overfitting and ensure the robustness of the SFM method.
4.6 ODE solvers and speed analysis
Table 2: Adaptive-step ODE solvers used in our experiments. All of them are variants of the explicit Runge–Kutta family.
Abbreviation Full Name Order
Heun(2) Adaptive Heun’s Method 2
Fehlberg(2) Runge–Kutta–Fehlberg Method 2
Bosh(3) Bogacki–Shampine Method 3
Dopri(5) Dormand–Prince Method 5
For evaluations, we follow the official ODE solver settings of each model. For speed analysis with adaptive-step ODE solvers, we use the odeint from torchdiffeq torchdiffeq with various solvers. The default relative and absolute tolerances ($\times1$0-7 and $\times1$0-9) are too strict, resulting in significantly longer inference time. Therefore, we adopt the settings used in StableTTS: $\times1$0-5 for both tolerances. As shown in Table 2, we employ Huen(2), Fehlberg(2), Bosh(3), and Dopri(5) as adaptive-step solvers, where the number in () denotes the order.
We adopt real-time factor (RTF) and number of function evaluations (NFE) as metrics for speed evaluation, where RTF measures the ratio between the ODE solver inference time and the corresponding audio duration, and NFE indicates how many times the solver queries the model. To reduce the influence of runtime variance, we choose 100 utterances from each model’s validation set and run each solver five times, reporting the average RTF. Note that the NFE is constant across runs.
5 Experimental results and analysis
Table 3: Partial objective evaluation results on validation sets for $\boldsymbol{\alpha}$ selection. Complete results are provided in Appendix D. The highest value for each metric is highlighted in bold.
[…]
Matcha-TTS (SFM) trained on LJ Speech
1.0 0.099 0.092 4.036 4.194 3.721 4.192 4.641 0.972
2.0 0.198 0.183 4.158 4.305 3.834 4.337 3.496 0.973
2.5 0.248 0.229 4.176 4.276 3.872 4.381 3.556 0.972
3.0 0.297 0.275 4.168 4.260 3.842 4.402 3.496 0.970
3.5 0.347 0.320 4.132 4.190 3.802 4.403 3.496 0.969
4.0 0.397 0.366 4.107 4.137 3.763 4.421 3.556 0.966
5.0 0.496 0.458 4.025 3.977 3.694 4.403 3.376 0.960
6.0 0.520 0.480 3.997 3.958 3.648 4.386 3.315 0.958
7.0 0.520 0.480 3.990 3.960 3.625 4.386 3.315 0.958
8.0 0.520 0.480 3.990 3.959 3.625 4.386 3.315 0.958
9.0 0.520 0.480 3.993 3.956 3.638 4.386 3.315 0.958
10.0 0.520 0.480 3.987 3.955 3.620 4.386 3.315 0.958
5.1 Evaluation results
Due to page limitations, Table 3 and Table 5 only provide partial results, and complete results are provided in Appendix D and F, respectively. Other minor analysis is provided in Appendix E.
SFM strength selection: In Table 3 and the tables in Appendix D, $\tilde{t}_{\boldsymbol{g}}$ and $\tilde{\sigma}_{\boldsymbol{g}}$ are computed according to Eq. 22, and the reported values represent their means over all utterances in the validation sets. From these tables, it is evident that using values of $\alpha>1.0$ significantly improves both pseudo-MOS scores and WER. The $\overline{\text{PMOS}}$, which serves as the selection criterion, tends to increase initially and then decrease as $\alpha$ grows. This observation suggests that the adaptively determined $t_{\boldsymbol{h}}$ during training is generally small. As a result, the intermediate state constructed at inference with $\alpha=1.0$ corresponds to an early position on the CondOT paths. When Gaussian noise $\boldsymbol{X}_{0}$ is added at this stage, the resulting signal-to-noise ratio is low, making it difficult for the flow decoder to extract sufficient information and weakening the guidance during early sampling steps.
According to Eq. (13), $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ is supervised by the linearly down-scaled $t_{\boldsymbol{h}}\boldsymbol{X}_{1}$. This allows us to apply the $\alpha$ during inference to linearly scale up $\tilde{\boldsymbol{X}}_{\boldsymbol{h}}$ in Eq. 22, thereby enhancing its guidance effect. However, due to estimation errors, increasing $\alpha$ also leads to a growing distance between $\tilde{\boldsymbol{X}}_{\boldsymbol{h}}$ and $\tilde{t}_{\boldsymbol{h}}\boldsymbol{X}_{1}$. Meanwhile, the incorporated $\boldsymbol{X}_{0}$ decreases and results in a more deterministic sampling process. These factors ultimately reduce generation quality.
Table 4: Evaluation results on test sets. * indicates statistically significant differences ($p<0.05$) compared with SFM models in subjective evaluations. The highest value for each metric is bolded.
[…]
Matcha-TTS trained on LJ Speech
Ground truth 4.380 3.964 4.241 3.566 1.000 +0.22+0.22
Reconstructed 4.085 3.739 4.208 3.472 0.993 +0.12+0.12
Baseline 4.186 3.692 4.282 3.308 0.971 0.48-0.48
Ablated 4.217 3.763 4.311 3.355 0.972 0.27-0.27
SFM ($\alpha$=2.5) 4.257 3.848 4.386 3.413 0.972 0.00
Matcha-TTS trained on VCTK
Ground truth 3.999 3.562 3.986 1.534 1.000 +0.16+0.16
Reconstructed 3.819 3.246 3.977 1.666 0.985 +0.08+0.08
Baseline 4.008 2.978 3.870 1.534 0.939 0.31-0.31*
Ablated 4.026 2.997 3.872 1.613 0.941 0.39-0.39*
SFM ($\alpha$=3.5) 4.106 3.105 3.898 0.952 0.937 0.00
StableTTS trained on VCTK
Ground truth 3.999 3.562 3.986 1.534 1.000 +0.48+0.48*
Reconstructed 3.360 2.908 3.855 1.719 0.972 +0.04+0.04
Ablated 3.328 2.958 3.929 1.798 0.932 0.34-0.34*
SFM ($\alpha$=3.0) 3.516 3.020 3.953 1.745 0.933 0.00
SFM-c ($\alpha$=4.5) 3.507 2.899 3.934 1.877 0.931 0.37-0.37*
CosyVoice trained on LibriTTS
Ground truth 4.136 3.262 4.345 3.180 1.000 +0.19+0.19 3.40
Reconstructed 3.942 3.126 4.336 3.146 0.993 0.24-0.24 2.82*
Baseline 4.191 3.303 4.481 3.513 0.932 0.21-0.21* 3.47
Ablated 4.183 3.369 4.487 3.578 0.932 0.14-0.14 3.58
SFM ($\alpha$=2.0) 4.194 3.480 4.541 3.810 0.931 0.00 3.67
SFM-t ($\alpha$=2.5) 4.132 3.336 4.547 3.987 0.914 0.09-0.09 2.66*
CosyVoice-DiT trained on LibriTTS
Ground truth 4.136 3.262 4.345 3.180 1.000 +0.23+0.23 3.31
Reconstructed 3.322 2.855 4.211 3.144 0.989 0.12-0.12 2.86*
Ablated 3.499 3.086 4.316 3.614 0.936 0.31-0.31* 3.15
SFM ($\alpha$=2.5) 3.751 3.171 4.502 3.598 0.932 0.00 3.21
SFM-c ($\alpha$=4.0) 3.752 3.156 4.496 3.634 0.929 0.06-0.06 3.10
Objective evaluations: As shown in Table 4, all SFM models outperform their corresponding baseline and ablated models regarding pseudo-MOS scores. However, the results for WER and SIM are more variable, with improvements observed in some cases but not in all. This suggests that there is still room to improve the alignment quality of the SFM method.
Subjective evaluations: As shown in Table 4, SFM models achieve better performance in both CMOS and SMOS tests compared to their corresponding baseline, ablated, and SFM variants. This demonstrates that the SFM method efficiently improves the naturalness of synthesized speech.
CosyVoice (SFM-t): When only speech tokens are input into the encoder of the flow module, CosyVoice (SFM-t) exhibits lower SIM and significantly worse SMOS performance. This is because the CosyVoice tokenizer is pre-trained with an ASR objective to extract semantic tokens, which contain limited speaker-specific information. Although speaker-related features are employed as flow conditions, they fail to guide the generated speech with sufficient speaker characteristics. This further highlights the importance of early-stage flow inference, as errors introduced at the beginning are difficult to correct in later sampling steps.
SFM-c: When using coarse mel-spectrograms as flow conditions, the adaptively determined $t_{\boldsymbol{h}}$ tends to converge to 0 in models with U-Net-based architectures, rendering the SFM-c variant inapplicable to these models. Models using DiT blocks, such as StableTTS (SFM-c) and CosyVoice-DiT (SFM-c), perform worse in subjective evaluations than their corresponding SFM models. These results suggest that, for our SFM method, using coarse mel-spectrograms as flow conditions not only fails to improve performance but can also degrade synthesis quality or invalidate the method.
5.2 Acceleration of adaptive-step ODE solvers
Table 5: Partial RTF and NFE results for adaptive-step ODE solvers. Complete results are provided in Appendix F. $\overline{\text{RTF}}$ denotes the mean RTF. Rate (%) denotes the relative speedup in terms of $\overline{\text{RTF}}$ compared to the ablated model.
System Heun(2) Fehlberg(2) Bosh(3) Dopri(5)
[…]
Matcha-TTS trained on LJ Speech
Baseline 0.407 -1.496 312.36 0.054 3.571 44.82 0.271 -0.370 223.95 0.142 2.069 120.10
Ablated 0.401 0.000 306.72 0.056 0.000 45.28 0.270 0.000 221.81 0.145 0.000 121.46
SFM ($\alpha$=1.0) 0.361 9.858 277.56 0.053 4.171 44.08 0.225 16.924 186.23 0.132 9.100 111.14
SFM ($\alpha$=2.0) 0.299 25.541 229.43 0.048 13.306 39.16 0.187 30.931 153.08 0.115 20.484 96.02
SFM ($\alpha$=3.0) 0.249 37.931 191.49 0.043 22.749 34.88 0.157 41.895 129.02 0.100 30.753 84.20
SFM ($\alpha$=4.0) 0.207 48.486 156.78 0.037 34.246 29.74 0.131 51.576 107.90 0.086 40.559 72.56
SFM ($\alpha$=5.0) 0.157 60.825 120.07 0.027 50.968 22.14 0.110 59.266 90.74 0.076 47.605 63.74
Table 5 and the tables in Appendix F report only the mean RTF for clarity, as the standard deviations across the five runs are all below 0.013. We also present the speedup rate in terms of RTF, measured relative to the ablated model. From these tables, we observe that increasing $\alpha$ improves the signal-to-noise ratio of the initial state during flow inference, which stabilizes the ODE solving process. As a result, fewer forward steps are required, and the overall inference time is significantly reduced. Notably, this acceleration effect is limited to adaptive-step solvers, as fixed-step solvers perform a predefined number of steps, and thus cannot leverage the improved stability of the initial stages to reduce inference cost.
6 Related works
Our proposed SFM method extends the idea of the shallow diffusion mechanism. To the best of our knowledge, while no prior work exactly matches our approach, several studies adopt similar strategies or pursue similar objectives. As discussed in Section 3.1, ReRFlow PeRFlow applies piecewise reflow to divided flow trajectories. PixelFlow also adopts piecewise flow for multi-scale resolution generation. In our case, we utilize piecewise flow to divide the CondOT paths into two segments and use only the last segment. In addition, shortcut models shortcut employ a self-consistency mechanism to construct shortcuts along the CondOT paths. Modifying flow matching modifyingflow samples from a Gaussian distribution centered at a coarse output instead of the standard normal distribution, and further adopts deterministic inference.
7 Limitations
This work applies only minimal modifications to the backbone TTS models and uses a simple SFM head to demonstrate the effectiveness of our proposed SFM method. Therefore, there is considerable room for improving the SFM framework and corresponding implementation. For example:
1. A more powerful SFM head could be used when the weak generator is unstable.
2. The flow conditions could be enhanced by incorporating text or speech token embeddings to better align the final outputs with the input semantic features.
3. When applying SFM on CosyVoice, we directly concatenate the speech tokens, speaker embeddings, and masked mel-spectrograms as the encoder input. This naive fusion strategy may negatively affect cross-modal alignment and calls for a more elaborate integration approach.
8 Conclusion
We introduce a novel Shallow Flow Matching (SFM) method for coarse-to-fine TTS models. SFM leverages coarse representations to construct intermediate states on the CondOT path, enabling flow-based inference to produce more stable generation, more natural synthetic speech, and faster inference when using adaptive-step ODE solvers.
SFM focuses on cases where the FM module serves as a refiner. Although it is validated on TTS tasks, the underlying framework and theoretical foundation are general. It also holds potential for other domains, such as speech denoising and enhancement, as well as image generation tasks like denoising and super-resolution. Further exploration of its applications to these tasks remains a promising direction for future work.
9 Acknowledgements
This research was supported by JST Moonshot Grant Number JPMJMS2237 and JST SPRING Grant Number JPMJSP2108.
Appendix A Theorem Proofs
Theorem 1.
For any random variable $\boldsymbol{x}_{m}\sim\mathcal{N}(t_{m}\boldsymbol{x}_{1},\sigma^{2}_{m}\boldsymbol{I})$, where $t_{m}\in[0,\infty)$ and $\sigma_{m}\in(0,\infty)$, we define a transformation that maps $\boldsymbol{x}_{m}$ onto the CondOT paths. The output distribution varies continuously with respect to $t_{m}$ and $\sigma_{m}$ under the Wasserstein-2 metric.
$\Delta=(1-\sigma_{\mathrm{min}})t_{m}+\sigma_{m},$ (23)
$\boldsymbol{x}_{\tau}$ $=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}}\boldsymbol{x}_{0}+\boldsymbol{x}_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}\boldsymbol{x}_{m},&\text{if }\Delta\geq 1,\end{cases}$ (24)
where $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, with corresponding $\tau=\min(t_{m},\frac{t_{m}}{\Delta})$ on the path.
Proof. Since $\boldsymbol{x}_{m}\sim\mathcal{N}(t_{m}\boldsymbol{x}_{1},\sigma^{2}_{m}\boldsymbol{I})$ and $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ are independent, as a linear combination of them, $\boldsymbol{x}_{\tau}$ also follows a Gaussian distribution with:
[…]
$\sigma(\boldsymbol{x}_{\tau})$ $=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}+\sigma^{2}_{m}}=1-(1-\sigma_{\mathrm{min}})\tau,&\text{if }\Delta<1,\\ \frac{\sigma_{m}}{\Delta}=1-(1-\sigma_{\mathrm{min}})\frac{t_{m}}{\Delta}=1-(1-\sigma_{\mathrm{min}})\tau,&\text{if }\Delta\geq 1,\end{cases}$ (27)
where $\boldsymbol{\mu}(\cdot)$ denotes the mean and $\sigma(\cdot)$ denotes the std.
Since $\boldsymbol{\mu}(\boldsymbol{x}_{\tau})$ and $\sigma(\boldsymbol{x}_{\tau})$ satisfy Eq. (3), $\boldsymbol{x}_{\tau}$ lies on the CondOT paths.
We denote the transformation as $T:[0,1]\times(0,\infty)\times\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Since $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and $\boldsymbol{x}_{1}$ is a sample from the dataset, the distribution of $\boldsymbol{x}_{\tau}$ depends only on $t_{m}$ and $\sigma_{m}$, conditioned on $\boldsymbol{x}_{1}$. Therefore, we denote the distribution of $\boldsymbol{x}_{\tau}$ as:
$\boldsymbol{x}_{\tau}\sim\mathcal{N}(\boldsymbol{\mu}(\boldsymbol{x}_{\tau}),\sigma(\boldsymbol{x}_{\tau})^{2}\boldsymbol{I})=T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}).$ (28)
To verify that $T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m})$ is continuous in the Wasserstein-2 metric with respect to $t_{m}$ and $\sigma_{m}$, we use formula:
$W_{2}(\mathcal{N}_{1}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1}),\mathcal{N}_{2}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2}))=\sqrt{\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+\mathrm{Tr}\left(\boldsymbol{\Sigma}_{1}+\boldsymbol{\Sigma}_{2}-2(\boldsymbol{\Sigma}_{1}^{1/2}\boldsymbol{\Sigma}_{2}\boldsymbol{\Sigma}_{1}^{1/2})^{1/2}\right)}.$ (29)
If our distributions are isotropic, we have:
$W_{2}(\mathcal{N}_{1}(\boldsymbol{\mu}_{1},\sigma_{1}^{2}\boldsymbol{I}),\mathcal{N}_{2}(\boldsymbol{\mu}_{2},\sigma_{2}^{2}\boldsymbol{I})=\sqrt{\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|^{2}+n(\sigma_{1}-\sigma_{2})^{2}},$ (30)
where $n$ is the dimension of the variables. We define $\delta>0$ for proof.
For $t_{m}$:
I. When $t_{m}<\frac{1-\sigma_{m}}{1-\sigma_{\mathrm{min}}}$, then $\Delta<1$:
$\lim_{\delta\to 0}W_{2}(T_{\boldsymbol{x}_{1}}(t_{m}-\delta,\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}))$ $=\lim_{\delta\to 0}\sqrt{\|\delta\boldsymbol{x}_{1}\|^{2}+n(\delta(1-\sigma_{\mathrm{min}}))^{2}}$ (31)
$=0,$ (32)
$\lim_{\delta\to 0}W_{2}(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m}+\delta,\sigma_{m}))$ $=\lim_{\delta\to 0}\sqrt{\|\delta\boldsymbol{x}_{1}\|^{2}+n(\delta(1-\sigma_{\mathrm{min}}))^{2}}$ (33)
$=0.$ (34)
II. When $t_{m}=\frac{1-\sigma_{m}}{1-\sigma_{\mathrm{min}}}$, then $\Delta=1$:
$\lim_{\delta\to 0}W_{2}(T_{\boldsymbol{x}_{1}}(t_{m}-\delta,\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}))$ $=\lim_{\delta\to 0}\sqrt{\|\delta\boldsymbol{x}_{1}\|^{2}+n(\delta(1-\sigma_{\mathrm{min}}))^{2}}$ (35)
$=0,$ (36)
$\lim_{\delta\to 0}W_{2}$ $(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m}+\delta,\sigma_{m}))$ (37)
$=\lim_{\delta\to 0}\sqrt{\|[t_{m}-\frac{t_{m}+\delta}{1+\delta(1-\sigma_{\mathrm{min}})}]\boldsymbol{x}_{1}\|^{2}+n(\sigma_{m}-\frac{\sigma_{m}}{1+\delta(1-\sigma_{\mathrm{min}})})^{2}}$ (38)
$=0.$ (39)
III. When $t_{m}>\frac{1-\sigma_{m}}{1-\sigma_{\mathrm{min}}}$, then $\Delta>1$:
$\lim_{\delta\to 0}W_{2}$ $(T_{\boldsymbol{x}_{1}}(t_{m}-\delta,\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}))$ (40)
$=\lim_{\delta\to 0}\sqrt{\|[\frac{t_{m}-\delta}{\Delta-\delta(1-\sigma_{\mathrm{min}})}-\frac{t_{m}}{\Delta}]\boldsymbol{x}_{1}\|^{2}+n(\frac{\sigma_{m}}{\Delta-\delta(1-\sigma_{\mathrm{min}})}-\frac{\sigma_{m}}{\Delta})^{2}}$ (41)
$=0,$ (42)
$\lim_{\delta\to 0}W_{2}$ $(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m}+\delta,\sigma_{m}))$ (43)
$=\lim_{\delta\to 0}\sqrt{\|[\frac{t_{m}}{\Delta}-\frac{t_{m}+\delta}{\Delta+\delta(1-\sigma_{\mathrm{min}})}]\boldsymbol{x}_{1}\|^{2}+n(\frac{\sigma_{m}}{\Delta}-\frac{\sigma_{m}}{\Delta+\delta(1-\sigma_{\mathrm{min}})})^{2}}$ (44)
$=0.$ (45)
For $\sigma_{m}$:
I. When $\sigma_{m}<1-(1-\sigma_{\mathrm{min}})t_{m}$, then $\Delta<1$:
$\lim_{\delta\to 0}W_{2}(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}-\delta),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}))$ $=\lim_{\delta\to 0}\sqrt{\|0\boldsymbol{x}_{1}\|^{2}+n(0))^{2}}$ (46)
$=0,$ (47)
$\lim_{\delta\to 0}W_{2}(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}+\delta))$ $=\lim_{\delta\to 0}\sqrt{\|0\boldsymbol{x}_{1}\|^{2}+n(0))^{2}}$ (48)
$=0.$ (49)
II. When $\sigma_{m}=1-(1-\sigma_{\mathrm{min}})t_{m}$, then $\Delta=1$:
$\lim_{\delta\to 0}W_{2}(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}-\delta),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}))$ $=\lim_{\delta\to 0}\sqrt{\|0\boldsymbol{x}_{1}\|^{2}+n(0))^{2}}$ (50)
$=0,$ (51)
$\lim_{\delta\to 0}W_{2}$ $(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}+\delta))$ (52)
$=\lim_{\delta\to 0}\sqrt{\|[t_{m}-\frac{t_{m}}{1+\delta}]\boldsymbol{x}_{1}\|^{2}+n(\sigma_{m}-\frac{\sigma_{m}+\delta}{1+\delta})^{2}}$ (53)
$=0.$ (54)
III. When $\sigma_{m}>1-(1-\sigma_{\mathrm{min}})t_{m}$, then $\Delta>1$:
$\lim_{\delta\to 0}W_{2}$ $(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}-\delta),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}))$ (55)
$=\lim_{\delta\to 0}\sqrt{\|[\frac{t_{m}}{\Delta-\delta}-\frac{t_{m}}{\Delta}]\boldsymbol{x}_{1}\|^{2}+n(\frac{\sigma_{m}-\delta}{\Delta-\delta}-\frac{\sigma_{m}}{\Delta})^{2}}$ (56)
$=0,$ (57)
$\lim_{\delta\to 0}W_{2}$ $(T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}),T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m}+\delta))$ (58)
$=\lim_{\delta\to 0}\sqrt{\|[\frac{t_{m}}{\Delta}-\frac{t_{m}}{\Delta+\delta}]\boldsymbol{x}_{1}\|^{2}+n(\frac{\sigma_{m}}{\Delta}-\frac{\sigma_{m}+\delta}{\Delta+\delta})^{2}}$ (59)
$=0.$ (60)
Thus, $T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m})$ is continuous with respect to $t_{m}$ and $\sigma_{m}$ in the Wasserstein-2 metric. ∎
Theorem 2.
For arbitrary intermediate states on the conditional OT (CondOT) paths:
$\boldsymbol{x}_{t_{m}}=(1-t_{m})\boldsymbol{x}_{0}+t_{m}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),\quad t_{m}\in(0,1),\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}),$ (61)
we can divide the paths into two segments at $t_{m}$ and represent the flow and VF using piecewise functions:
$\boldsymbol{x}_{t}$ $=\begin{cases}(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}},&\text{if }t<t_{m},\\ (1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),&\text{if }t\geq t_{m},\end{cases}$ (62)
$\boldsymbol{u}_{t}$ $=\begin{cases}\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0}),&\text{if }t<t_{m},\\ \frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}}),&\text{if }t\geq t_{m}.\end{cases}$ (63)
In the first segment ($t<t_{m}$) of the two-segment piecewise flow, the paths start from $\boldsymbol{x}_{0}$ with $\boldsymbol{x}_{t_{m}}$ as the target. In the second segment ($t\geq t_{m}$), the paths start from $\boldsymbol{x}_{t_{m}}$ and move towards the target $(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})$.
Proof. Substituting Eq. (61) into Eq. (62) and Eq. (63), we derive the following:
I. When $t<t_{m}$}:
$\boldsymbol{x}_{t}$ $=(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}}$ (64)
$=(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}(1-t_{m})\boldsymbol{x}_{0}+t(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})$ (65)
$=(1-t)\boldsymbol{x}_{0}+t(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),$ (66)
$\boldsymbol{u}_{t}$ $=\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0})$ (67)
$=\frac{1}{t_{m}}(t_{m}\boldsymbol{x}_{0}+t_{m}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}))$ (68)
$=(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})-\boldsymbol{x}_{0}.$ (69)
II. When $t\geq t_{m}$}:
$\boldsymbol{x}_{t}$ $=(1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})$ (70)
$=(1-t)\boldsymbol{x}_{0}+\frac{1-t}{1-t_{m}}t_{m}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})$ (71)
$=(1-t)\boldsymbol{x}_{0}+t(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),$ (72)
$\boldsymbol{u}_{t}$ $=\frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}})$ (73)
$=\frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})-\boldsymbol{x}_{0}-\frac{t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})$ (74)
$=(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0})-\boldsymbol{x}_{0}.$ (75)
Therefore, the piecewise flow and VF above satisfy Eq. (4). ∎
Appendix B SFM head
Figure 2: SFM head architecture.
As shown in Fig. 2, we use a simple and lightweight SFM head, whose architecture is derived from the duration predictor of VITS [^4] https://github.com/jaywalnut310/vits/blob/2e561ba58618d021b5b8323d3765880f7e0ecfdb/models.py#L98 and Matcha-TTS. [^5] https://github.com/shivammehta25/Matcha-TTS/blob/108906c603fad5055f2649b3fd71d2bbdf222eac/matcha/models/components/text_encoder.py#L70 The kernel size and padding of the two Conv1d layers are set to 3 and 1, respectively.
The output of the SFM head can be split into $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$, $\hat{t}_{\boldsymbol{h}}$, and $\log\hat{\sigma}^{2}_{\boldsymbol{h}}$, with shapes [batch, mel, sequence], [batch, 1, sequence], and [batch, 1, sequence], respectively. Since $\hat{t}_{\boldsymbol{h}}>0$, we apply a Sigmoid activation to it and compute the mean over the frames axis, resulting in an output of shape [batch, 1]. As described in Appendix C, the SFM head predicts $\log\hat{\sigma}^{2}_{\boldsymbol{h}}$ instead of $\hat{\sigma}^{2}_{\boldsymbol{h}}$. $\log\hat{\sigma}^{2}_{\boldsymbol{h}}$ is also averaged across frames to produce an output of shape [batch, 1].
Appendix C Algorithms
In the practical implementation, we apply several detailed modifications and tricks:
1. The SFM-head $\boldsymbol{h_{\psi}}$ predicts $\log\hat{\sigma}^{2}_{\boldsymbol{h}}$ instead of $\hat{\sigma}^{2}_{\boldsymbol{h}}$ directly to ensure numerical stability and positive outputs.
2. Due to the $\mathcal{L}_{\mathrm{coarse}}$, the $t_{\boldsymbol{h}}$ obtained from $\boldsymbol{X}_{\boldsymbol{h}}$ are generally larger than 0 during training. Therefore, we remove the constraint $t_{\boldsymbol{h}}\geq 0$ in Equation 13.
3. We also modify the placement of the $\mathcal{L}_{\mu}$ for implementation convenience. Compared to the formulation in Equation 13, this modification results in a scaled version of the original loss by a constant factor of $\frac{1}{\Delta^{2}}$.
Algorithm 1 Training procedure of SFM
Input: the training set $(\mathcal{X},\mathcal{C})$.
1:repeat
2:  Sample $(\boldsymbol{X}_{1},\boldsymbol{C})$ from $(\mathcal{X},\mathcal{C})$;
3:  Generate coarse representations
$\boldsymbol{H_{g}},\boldsymbol{X_{g}}\leftarrow\boldsymbol{g_{\omega}}(\boldsymbol{C})$
$\boldsymbol{X}_{\boldsymbol{h}},\hat{t}_{\boldsymbol{h}},\log\hat{\sigma}^{2}_{\boldsymbol{h}}\leftarrow\boldsymbol{h_{\psi}}(\boldsymbol{H}_{\boldsymbol{g}})$
4:  Compute coarse loss
[…]
5:  Project onto CondOT paths
[…]
6:  Construct the intermediate state
$\Delta\leftarrow\max((1-\sigma_{\mathrm{min}})t_{\boldsymbol{h}}+\sigma_{\boldsymbol{h}},1)$
$\boldsymbol{X}_{\boldsymbol{h}}\leftarrow\frac{1}{\Delta}{\boldsymbol{X}_{\boldsymbol{h}}},\quad t_{\boldsymbol{h}}\leftarrow\frac{1}{\Delta}{t_{\boldsymbol{h}}},\quad\sigma^{2}_{\boldsymbol{h}}\leftarrow\frac{1}{\Delta^{2}}{\sigma^{2}_{\boldsymbol{h}}}$
$\boldsymbol{X}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$
$\boldsymbol{X}_{\boldsymbol{h}}\leftarrow\sqrt{\max((1-(1-\sigma_{\mathrm{min}})t_{\boldsymbol{h}})^{2}-\sigma_{\boldsymbol{h}}^{2},0)}\boldsymbol{X}_{0}+\boldsymbol{X}_{\boldsymbol{h}}$
[…]
7:  Apply FM
$t\sim\mathcal{U}[0,1],\quad t\leftarrow\mathcal{S}(t)$
$\boldsymbol{X}_{t}\leftarrow(1-t)\boldsymbol{X}_{\boldsymbol{h}}+t(\boldsymbol{X}_{1}+\sigma_{\mathrm{min}}\boldsymbol{X}_{0})$
$\boldsymbol{U}_{t}\leftarrow\frac{1}{1-t_{\boldsymbol{h}}}(\boldsymbol{X}_{1}+\sigma_{\mathrm{min}}\boldsymbol{X}_{0}-\boldsymbol{X}_{\boldsymbol{h}})$
$t\leftarrow(1-t_{\boldsymbol{h}})t+t_{\boldsymbol{h}}$
[…]
8:  Apply gradient descent to minimize $\mathcal{L}_{\mathrm{SFM}}$
$\mathcal{L}_{\mathrm{SFM}}=\mathcal{L}_{\mathrm{coarse}}+\mathcal{L}_{t}+\mathcal{L}_{\sigma}+\mathcal{L}_{\mu}+\mathcal{L}_{\mathrm{CFM}}$
9:until convergence
Algorithm 2 Inference procedure of SFM
Input: Condition $\boldsymbol{C}$.
1:Generate coarse representations
$\boldsymbol{H_{g}},\boldsymbol{X_{g}}\leftarrow\boldsymbol{g_{\omega}}(\boldsymbol{C})$
$\boldsymbol{X}_{\boldsymbol{h}},t_{\boldsymbol{h}},\log\sigma^{2}_{\boldsymbol{h}}\leftarrow\boldsymbol{h_{\psi}}(\boldsymbol{H}_{\boldsymbol{g}})$
$\sigma^{2}_{\boldsymbol{h}}\leftarrow\exp(\log\sigma^{2}_{\boldsymbol{h}})$
2:Construct the intermediate state with SFM strength
$\Delta\leftarrow\max(\alpha((1-\sigma_{\mathrm{min}})t_{\boldsymbol{h}}+\sigma_{\boldsymbol{h}}),1)$
$\boldsymbol{X}_{\boldsymbol{h}}\leftarrow\frac{\alpha}{\Delta}{\boldsymbol{X}_{\boldsymbol{h}}},\quad t_{\boldsymbol{h}}\leftarrow\frac{\alpha}{\Delta}{t_{\boldsymbol{h}}},\quad\sigma^{2}_{\boldsymbol{h}}\leftarrow\frac{\alpha^{2}}{\Delta^{2}}{\sigma^{2}_{\boldsymbol{h}}}$
$\boldsymbol{X}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$
$\boldsymbol{X}_{\boldsymbol{h}}\leftarrow\sqrt{\max((1-(1-\sigma_{\mathrm{min}})t_{\boldsymbol{h}})^{2}-\sigma_{\boldsymbol{h}}^{2},0)}\boldsymbol{X}_{0}+\boldsymbol{X}_{\boldsymbol{h}},\quad$
3:Use an ODE solver to solve the integral
$\boldsymbol{X}_{\mathrm{pred}}=\boldsymbol{X}_{\boldsymbol{h}}+\int_{t_{\boldsymbol{h}}}^{1}\boldsymbol{v_{\theta}}(\boldsymbol{X}_{t},t)\,dt$
Appendix D Complete results of $\boldsymbol{\alpha}$ selection
We provide complete results in optimal $\alpha$ selection for all models in this section, including Table 6, Table 7, Table 8, and Table 9.
Table 6: Objective evaluation results on validation sets for $\alpha$ selection (Matcha-TTS).
[…]
Matcha-TTS (SFM) trained on LJ Speech
1.0 0.099 0.092 4.036 4.194 3.721 4.192 4.641 0.972
2.0 0.198 0.183 4.158 4.305 3.834 4.337 3.496 0.973
2.5 0.248 0.229 4.176 4.276 3.872 4.381 3.556 0.972
3.0 0.297 0.275 4.168 4.260 3.842 4.402 3.496 0.970
3.5 0.347 0.320 4.132 4.190 3.802 4.403 3.496 0.969
4.0 0.397 0.366 4.107 4.137 3.763 4.421 3.556 0.966
5.0 0.496 0.458 4.025 3.977 3.694 4.403 3.376 0.960
6.0 0.520 0.480 3.997 3.958 3.648 4.386 3.315 0.958
7.0 0.520 0.480 3.990 3.960 3.625 4.386 3.315 0.958
8.0 0.520 0.480 3.990 3.959 3.625 4.386 3.315 0.958
9.0 0.520 0.480 3.993 3.956 3.638 4.386 3.315 0.958
10.0 0.520 0.480 3.987 3.955 3.620 4.386 3.315 0.958
Matcha-TTS (SFM) trained on VCTK
1.0 0.100 0.078 3.462 3.876 2.723 3.787 4.898 0.923
2.0 0.201 0.155 3.647 4.061 2.988 3.891 2.177 0.940
2.5 0.251 0.194 3.652 4.045 3.007 3.904 2.041 0.940
3.0 0.301 0.233 3.678 4.079 3.041 3.914 1.497 0.939
3.5 0.351 0.272 3.679 4.074 3.053 3.910 1.224 0.936
4.0 0.401 0.311 3.660 4.079 3.001 3.900 1.088 0.934
5.0 0.502 0.389 3.615 4.059 2.916 3.869 1.088 0.929
6.0 0.564 0.436 3.555 4.084 2.757 3.824 1.224 0.923
7.0 0.564 0.436 3.554 4.084 2.752 3.824 1.224 0.923
8.0 0.564 0.436 3.563 4.084 2.782 3.824 1.224 0.923
9.0 0.564 0.436 3.559 4.082 2.770 3.825 1.224 0.923
10.0 0.564 0.436 3.546 4.083 2.730 3.824 1.224 0.923
Table 7: Objective evaluation results on validation sets for $\alpha$ selection (StableTTS).
[…]
StableTTS (SFM) trained on VCTK
1.0 0.153 0.098 3.281 3.184 2.782 3.877 8.707 0.910
2.0 0.306 0.197 3.441 3.403 2.981 3.938 2.041 0.931
2.5 0.382 0.246 3.476 3.469 3.013 3.945 1.361 0.933
3.0 0.459 0.295 3.486 3.522 3.001 3.936 1.224 0.933
3.5 0.535 0.345 3.475 3.543 2.959 3.923 1.224 0.933
4.0 0.603 0.388 3.437 3.547 2.857 3.906 1.361 0.933
5.0 0.609 0.391 3.435 3.560 2.829 3.914 1.361 0.933
6.0 0.609 0.391 3.434 3.561 2.825 3.915 1.361 0.933
7.0 0.609 0.391 3.431 3.558 2.819 3.915 1.361 0.933
8.0 0.609 0.391 3.434 3.559 2.828 3.915 1.361 0.933
9.0 0.609 0.391 3.435 3.560 2.832 3.914 1.361 0.933
10.0 0.609 0.391 3.433 3.560 2.824 3.914 1.361 0.933
StableTTS (SFM-c) trained on VCTK
1.0 0.111 0.067 3.261 3.165 2.759 3.859 1.224 0.918
2.0 0.222 0.134 3.313 3.260 2.787 3.893 1.633 0.922
3.0 0.333 0.201 3.403 3.375 2.913 3.923 1.497 0.927
3.5 0.388 0.235 3.424 3.419 2.916 3.936 1.361 0.928
4.0 0.444 0.268 3.427 3.462 2.889 3.931 1.088 0.930
4.5 0.499 0.302 3.449 3.495 2.924 3.929 1.088 0.931
5.0 0.555 0.335 3.408 3.510 2.799 3.913 1.361 0.931
6.0 0.624 0.376 3.377 3.553 2.672 3.906 1.224 0.934
7.0 0.624 0.376 3.389 3.554 2.707 3.906 1.224 0.934
8.0 0.624 0.376 3.398 3.554 2.734 3.907 1.224 0.934
9.0 0.624 0.376 3.387 3.553 2.702 3.907 1.224 0.934
10.0 0.624 0.376 3.378 3.554 2.674 3.906 1.224 0.934
Table 8: Objective evaluation results on validation sets for $\alpha$ selection (CosyVoice).
[…]
CosyVoice (SFM) trained on LibriTTS
1.0 0.152 0.126 3.721 3.729 3.102 4.333 8.902 0.923
1.5 0.228 0.189 4.019 4.113 3.449 4.494 4.810 0.932
2.0 0.303 0.253 4.087 4.180 3.530 4.553 4.499 0.931
2.5 0.379 0.316 4.080 4.165 3.516 4.558 4.475 0.928
3.0 0.455 0.379 4.019 4.119 3.409 4.527 4.523 0.922
4.0 0.546 0.454 3.917 4.075 3.235 4.440 4.427 0.916
5.0 0.546 0.454 3.911 4.081 3.209 4.444 4.475 0.916
6.0 0.546 0.454 3.904 4.070 3.199 4.443 4.523 0.916
7.0 0.546 0.454 3.905 4.070 3.201 4.442 4.475 0.916
8.0 0.546 0.454 3.912 4.075 3.220 4.440 4.427 0.916
9.0 0.546 0.454 3.919 4.077 3.239 4.441 4.475 0.916
10.0 0.546 0.454 3.911 4.081 3.207 4.444 4.475 0.916
CosyVoice (SFM-t) trained on LibriTTS
1.0 0.088 0.152 3.567 3.501 2.957 4.242 14.932 0.917
1.5 0.132 0.228 3.872 3.932 3.267 4.417 6.102 0.928
2.0 0.176 0.304 4.000 4.089 3.392 4.520 4.882 0.924
2.5 0.219 0.380 4.005 4.078 3.395 4.543 4.834 0.911
3.0 0.263 0.456 3.877 3.940 3.213 4.477 4.355 0.896
4.0 0.349 0.605 3.395 3.498 2.599 4.087 4.355 0.868
5.0 0.364 0.635 3.251 3.375 2.421 3.955 4.379 0.865
6.0 0.364 0.636 3.260 3.376 2.441 3.963 4.331 0.865
7.0 0.364 0.636 3.261 3.380 2.450 3.953 4.355 0.864
8.0 0.364 0.636 3.271 3.379 2.470 3.963 4.403 0.865
9.0 0.364 0.636 3.271 3.377 2.478 3.958 4.427 0.865
10.0 0.364 0.636 3.255 3.375 2.433 3.955 4.379 0.865
Table 9: Objective evaluation results on validation sets for $\alpha$ selection (CosyVoice-DiT).
[…]
CosyVoice-DiT (SFM) trained on LibriTTS
1.0 0.143 0.120 3.405 3.077 2.880 4.257 8.423 0.924
2.0 0.286 0.239 3.750 3.569 3.189 4.493 4.499 0.934
2.5 0.358 0.299 3.823 3.694 3.247 4.529 4.331 0.933
3.0 0.429 0.359 3.810 3.720 3.184 4.528 4.212 0.929
3.5 0.501 0.419 3.780 3.758 3.089 4.493 4.068 0.925
4.0 0.545 0.455 3.721 3.750 2.968 4.444 4.092 0.922
5.0 0.545 0.455 3.722 3.750 2.972 4.444 4.092 0.922
6.0 0.545 0.455 3.720 3.749 2.967 4.444 4.092 0.922
7.0 0.545 0.455 3.723 3.751 2.973 4.445 4.092 0.922
8.0 0.545 0.455 3.718 3.750 2.959 4.444 4.092 0.922
9.0 0.545 0.455 3.724 3.750 2.979 4.444 4.092 0.922
10.0 0.545 0.455 3.719 3.750 2.964 4.444 4.092 0.922
CosyVoice-DiT (SFM-c) trained on LibriTTS
1.0 0.106 0.082 3.419 3.154 2.943 4.161 4.666 0.933
2.0 0.212 0.164 3.655 3.408 3.146 4.409 4.379 0.937
3.0 0.317 0.246 3.766 3.573 3.217 4.509 4.307 0.936
3.5 0.370 0.287 3.799 3.640 3.230 4.526 4.283 0.934
4.0 0.423 0.329 3.815 3.696 3.227 4.522 4.283 0.933
4.5 0.476 0.370 3.814 3.739 3.180 4.521 4.331 0.931
5.0 0.529 0.411 3.794 3.789 3.080 4.513 4.164 0.929
6.0 0.563 0.437 3.780 3.795 3.056 4.488 4.116 0.927
7.0 0.563 0.437 3.767 3.794 3.021 4.488 4.140 0.927
8.0 0.563 0.437 3.770 3.795 3.029 4.488 4.140 0.927
9.0 0.563 0.437 3.767 3.792 3.019 4.488 4.116 0.927
10.0 0.563 0.437 3.776 3.794 3.047 4.488 4.164 0.927
Appendix E Minor analysis of evaluation results
Interestingly, the Matcha-TTS (SFM) model trained on VCTK achieves a much lower WER than the ground-truth and vocoder-reconstructed speech. Since VCTK is a reading-style corpus, we observe that speakers occasionally exhibit disfluencies such as stammering or slurring. The synthesized speech, being generated from clean transcripts, avoids such inconsistencies.
For models trained on LibriTTS, vocoder-reconstructed speech performs significantly worse in SMOS tests. Since we adopt cross-sentence evaluations [8, 7], the prompt and target utterances are from the same speaker but are not necessarily adjacent, often resulting in prosodic mismatches. In contrast, CosyVoice LLM can imitate the prompt’s prosody, and the generated speech tokens align well with the prompt. Ground truth still performs competitively, likely due to its superior speech quality.
Reconstructed speech also performs poorly in CMOS tests. As LibriTTS is from audiobooks, which often feature expressive and emotional delivery. We observe that listeners tend to perceive such utterances as unnatural when heard in isolation. While flow matching-based models are capable of generating high-quality speech, prosody becomes the dominant factor in listener ratings. Ground-truth speech can outperform synthesized speech in CMOS due to its better sound quality.
Appendix F RTF and NFE with using adaptive-step ODE solvers
In this section, we provide not only the RTFs for ODE solver inference in Table 10, but also the RTFs for the full model inference in Table 11. These comparisons reveal that the ODE solver inference dominates the overall inference time. When using adaptive-step ODE solvers, the inference acceleration of our SFM methods is significant.
Table 10: RTF and NFE results for adaptive-step ODE solvers (solver inference). $\overline{\text{RTF}}$ denotes the mean of RTF. Rate (%) denotes the relative speedup in terms of $\overline{\text{RTF}}$ compared to the ablated model.
System Heun(2) Fehlberg(2) Bosh(3) Dopri(5)
[…]
Matcha-TTS trained on LJ Speech
Baseline 0.407 -1.496 312.36 0.054 3.571 44.82 0.271 -0.370 223.95 0.142 2.069 120.10
Ablated 0.401 0.000 306.72 0.056 0.000 45.28 0.270 0.000 221.81 0.145 0.000 121.46
SFM ($\alpha$=1.0) 0.361 9.858 277.56 0.053 4.171 44.08 0.225 16.924 186.23 0.132 9.100 111.14
SFM ($\alpha$=2.0) 0.299 25.541 229.43 0.048 13.306 39.16 0.187 30.931 153.08 0.115 20.484 96.02
SFM ($\alpha$=3.0) 0.249 37.931 191.49 0.043 22.749 34.88 0.157 41.895 129.02 0.100 30.753 84.20
SFM ($\alpha$=4.0) 0.207 48.486 156.78 0.037 34.246 29.74 0.131 51.576 107.90 0.086 40.559 72.56
SFM ($\alpha$=5.0) 0.157 60.825 120.07 0.027 50.968 22.14 0.110 59.266 90.74 0.076 47.605 63.74
Matcha-TTS trained on VCTK
Baseline 0.972 -2.101 318.42 0.139 -6.107 49.00 0.681 1.304 243.40 0.365 1.617 133.92
Ablated 0.952 0.000 313.18 0.131 0.000 45.52 0.690 0.000 245.69 0.371 0.000 134.76
SFM ($\alpha$=1.0) 0.950 0.210 312.26 0.123 6.269 42.60 0.690 0.030 244.43 0.371 0.111 134.00
SFM ($\alpha$=2.0) 0.771 18.998 253.66 0.111 15.463 38.28 0.558 19.071 196.58 0.318 14.356 114.08
SFM ($\alpha$=3.0) 0.626 34.271 205.49 0.100 23.635 34.50 0.436 36.844 153.86 0.265 28.626 95.48
SFM ($\alpha$=4.0) 0.519 45.449 170.63 0.093 29.503 31.96 0.350 49.257 123.80 0.231 37.846 83.48
SFM ($\alpha$=5.0) 0.407 57.242 133.78 0.073 44.413 25.10 0.279 59.547 99.05 0.195 47.454 70.28
StableTTS trained on VCTK
Ablated 2.153 0.000 712.06 0.394 0.000 133.96 1.270 0.000 434.74 0.998 0.000 346.12
SFM ($\alpha$=1.0) 1.212 43.694 397.26 0.261 33.707 88.36 0.790 37.823 269.32 0.690 30.849 232.60
SFM ($\alpha$=2.0) 0.982 54.386 316.00 0.232 41.301 76.48 0.697 45.123 231.82 0.613 38.593 208.72
SFM ($\alpha$=3.0) 0.789 63.346 254.74 0.205 48.081 67.44 0.611 51.930 204.64 0.564 43.473 190.60
SFM ($\alpha$=4.0) 0.592 72.497 192.34 0.169 57.149 53.84 0.542 57.305 181.72 0.489 51.003 166.12
SFM ($\alpha$=5.0) 0.586 72.792 189.30 0.173 56.096 53.64 0.538 57.674 178.90 0.505 49.377 169.48
CosyVoice trained on LibriTTS
Baseline 2.416 4.581 390.06 0.358 2.981 58.62 1.894 12.920 327.25 1.294 7.373 215.06
Ablated 2.532 0.000 395.25 0.369 0.000 58.91 2.175 0.000 346.90 1.397 0.000 223.37
SFM ($\alpha$=1.0) 1.745 31.093 275.92 0.283 23.449 45.68 1.454 33.143 232.67 1.061 24.069 170.57
SFM ($\alpha$=2.0) 1.426 43.667 218.17 0.250 32.196 39.63 1.211 44.326 192.97 0.917 34.395 146.30
SFM ($\alpha$=3.0) 1.110 56.182 173.28 0.219 40.682 34.85 1.001 53.952 158.57 0.787 43.679 126.53
SFM ($\alpha$=4.0) 0.963 61.967 149.95 0.186 49.642 29.98 0.887 59.217 142.15 0.720 48.469 116.09
SFM ($\alpha$=5.0) 0.960 62.096 149.90 0.185 49.875 29.96 0.888 59.190 141.79 0.727 47.990 116.15
CosyVoice-DiT trained on LibriTTS
Ablated 1.170 0.000 834.46 0.188 0.000 137.56 0.699 0.000 514.30 0.539 0.000 403.36
SFM ($\alpha$=1.0) 0.621 46.934 421.90 0.125 33.704 86.84 0.396 43.324 277.72 0.337 37.591 234.70
SFM ($\alpha$=2.0) 0.503 56.980 335.96 0.108 42.475 75.28 0.336 51.894 237.88 0.286 46.910 207.94
SFM ($\alpha$=3.0) 0.391 66.558 267.38 0.092 51.292 63.82 0.295 57.811 207.58 0.257 52.252 184.24
SFM ($\alpha$=4.0) 0.300 74.320 205.89 0.073 61.121 51.72 0.259 62.958 182.17 0.229 57.595 163.48
SFM ($\alpha$=5.0) 0.302 74.212 205.67 0.074 60.825 51.74 0.258 63.061 182.23 0.229 57.523 163.06
Table 11: RTF results when using adaptive-step ODE solvers (model inference). $\overline{\text{RTF}}$ denotes the mean of RTF. Rate (%) denotes the relative speedup compared to the ablated model.
System Heun(2) Fehlberg(2) Bosh(3) Dopri(5)
[…]
Matcha-TTS trained on LJ Speech
Baseline 0.409 -1.489 0.055 3.509 0.273 -0.368 0.143 2.055
Ablated 0.403 0.000 0.057 0.000 0.272 0.000 0.146 0.000
SFM ($\alpha$=1.0) 0.363 9.809 0.055 4.017 0.226 16.814 0.133 8.976
SFM ($\alpha$=2.0) 0.300 25.430 0.050 12.866 0.188 30.735 0.117 20.230
SFM ($\alpha$=3.0) 0.251 37.773 0.045 22.063 0.159 41.639 0.102 30.399
SFM ($\alpha$=4.0) 0.208 48.282 0.038 33.261 0.133 51.268 0.088 40.104
SFM ($\alpha$=5.0) 0.159 60.578 0.029 49.541 0.112 58.914 0.077 47.076
Matcha-TTS trained on VCTK
Baseline 0.975 -2.094 0.142 -5.185 0.684 1.441 0.369 1.600
Ablated 0.955 0.000 0.135 0.000 0.694 0.000 0.375 0.000
SFM ($\alpha$=1.0) 0.953 0.211 0.127 6.111 0.693 0.036 0.374 0.116
SFM ($\alpha$=2.0) 0.775 18.925 0.115 15.046 0.562 18.971 0.321 14.213
SFM ($\alpha$=3.0) 0.629 34.141 0.104 23.000 0.439 36.651 0.268 28.349
SFM ($\alpha$=4.0) 0.523 45.274 0.096 28.711 0.354 48.998 0.234 37.481
SFM ($\alpha$=5.0) 0.411 57.021 0.077 43.209 0.283 59.235 0.199 46.993
StableTTS trained on VCTK
Ablated 2.157 0.000 0.398 0.000 1.274 0.000 1.001 0.000
SFM ($\alpha$=1.0) 1.216 43.622 0.265 33.396 0.793 37.716 0.694 30.730
SFM ($\alpha$=2.0) 0.986 54.294 0.235 40.912 0.701 44.992 0.616 38.447
SFM ($\alpha$=3.0) 0.793 63.242 0.208 47.643 0.614 51.781 0.568 43.315
SFM ($\alpha$=4.0) 0.596 72.380 0.172 56.640 0.546 57.146 0.492 50.821
SFM ($\alpha$=5.0) 0.589 72.674 0.177 55.598 0.541 57.513 0.509 49.200
CosyVoice trained on LibriTTS
Baseline 2.417 4.580 0.359 2.973 1.895 12.914 1.295 7.368
Ablated 2.533 0.000 0.370 0.000 2.176 0.000 1.398 0.000
SFM ($\alpha$=1.0) 1.746 31.076 0.284 23.353 1.455 33.124 1.062 24.035
SFM ($\alpha$=2.0) 1.428 43.644 0.252 32.070 1.212 44.301 0.918 34.360
SFM ($\alpha$=3.0) 1.111 56.155 0.220 40.536 1.003 53.922 0.788 43.638
SFM ($\alpha$=4.0) 0.964 61.937 0.187 49.471 0.888 59.184 0.721 48.424
SFM ($\alpha$=5.0) 0.961 62.066 0.186 49.705 0.889 59.157 0.728 47.946
CosyVoice-DiT trained on LibriTTS
Ablated 1.171 0.000 0.190 0.000 0.700 0.000 0.541 0.000
SFM ($\alpha$=1.0) 0.622 46.870 0.126 33.406 0.397 43.221 0.338 37.472
SFM ($\alpha$=2.0) 0.505 56.903 0.110 42.113 0.338 51.779 0.288 46.775
SFM ($\alpha$=3.0) 0.393 66.472 0.093 50.872 0.296 57.683 0.259 52.102
SFM ($\alpha$=4.0) 0.302 74.225 0.075 60.633 0.260 62.820 0.230 57.432
SFM ($\alpha$=5.0) 0.303 74.116 0.075 60.336 0.260 62.924 0.230 57.360